OpenStudy (anonymous):

How do I factor 36x^2-100y^2

OpenStudy (anonymous):

\[36x^2-100y^2\]

OpenStudy (mathmagician):

(6x-10y)(6x+10y)

OpenStudy (anonymous):

its a difference of two squares \[36x^2 - 100y^2 = (6x + 10y)(6x - 10y)\]

OpenStudy (anonymous):

when you multiply out, the (6x*10y) terms cancel each other

OpenStudy (anonymous):

that is why it works

OpenStudy (anonymous):

(6x)^2 - (10y)^2 Now We know that \[a ^{2} - b ^{2} = (a+b)(a-b)\] Therefore we get \[(6x+10y)(6x-10y)\] @eigenschmeigen got it before me :)

OpenStudy (anonymous):

(6x)^2-(10y)^2 a^2-b^2=(a+b)(a-b) Applying this identity : =(6x+10y)(6x-10y) these are the factors

OpenStudy (anonymous):

and @mathmagician before me :)

OpenStudy (anonymous):

Sorry. Did nt see @mathmagician :)

OpenStudy (anonymous):

Thanks guys. I originally got what you all got too but when I checked my answer online, it gave me: 4(3x-5y)(3x+5y) so I thought I did something wrong.

OpenStudy (anonymous):

another example would be \[25x^2 - 49y^2\] since \[25x^2 = (5x)^2 \text{ and } 49y^2 = (7y)^2\] \[25x^2 - 49y^2 = (5x-7y)(5x+7y)\] ah ok so it just wanted you to take out four as a factor as well